# RLeXplore: Accelerating Research in Intrinsically-Motivated Reinforcement Learning

### v1

@@ Appendix D Learning Curves
Figure 8: Learning curves of the baselines and RQ 1 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
@@
Figure 9: Learning curves of the RQ 2 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
@@
Figure 10: Learning curves of the RQ 3 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
@@
Figure 11: Learning curves of the RQ 4 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
@@
Figure 12: Learning curves of the RQ 5 on SuperMarioBros-1-1-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 13: Learning curves of RQ 6 on Procgen-1MazeHard. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 14: Learning curves of RQ 6 on Procgen-AllMazeHard. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
@@
Figure 15: Learning curves of RQ 7 on SuperMarioBrosRandomStages-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 16: Learning curves of RQ 8 (global+episodic exploration) on SuperMarioBros-1-1-v3 and SuperMarioBrosRandomStages-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
@@
Figure 17: Learning curves of RQ 8 (global+global exploration) on SuperMarioBros-1-1-v3 and SuperMarioBrosRandomStages-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Appendix E On-Policy RL Algorithms and Discrete Control Tasks
@@
In this section, we demonstrate the combination of RLeXplore and on-policy RL algorithms and its effectiveness on discrete control tasks. Specifically, we couple the PPO algorithm and intrinsic rewards, and evalulate their performance on Montezuma’s Revenge, a hard exploration task from the Atari benchmark [36]. Meanwhile, we utilize the PPO implementation of CleanRL [34] to show the adaptability of RLeXplore. Table 5 illustrates the training hyperparameters used for the experiments.
Table 5: Training hyperparameters for Montezuma’s Revenge.
@@
Appendix F Off-Policy RL Algorithms and Continuous Control Tasks
To showcase the generality of RLeXplore, we run additional experiments in settings different from the ones in the main paper. Concretely, we couple intrinsic rewards with soft actor-critic (SAC) [39], an off-policy RL algorithm, and test their performance in Ant-UMaze, a continuous control task with sparse rewards. Table 6 illustrates the training hyperparameters used for the experiments. We show the performance of Disagreement, RND, ICM and vanilla SAC in Figure 19. The results indicate that intrinsically-motivated agents are able to navigate the maze more efficiently, finding the goals more often than the vanilla agents that can only learn from the sparse task rewards.
We only use 3 intrinsic rewards with SAC because of the episodic nature of the other intrinsic reward methods. For example, the episodic memory in RIDE, PseudoCounts, NGU; and the episodic ellipsoid in E3B require the replay buffer to sample entire episodes instead of random rollouts. We aim to implement this logic in the future in our RLeXplore codebase.
@@ Appendix F Off-Policy RL Algorithms and Continuous Control Tasks
Policy learning rate $3\cdot 10^{-4}$

### v2

@@ Appendix E Learning Curves
Figure 10: Learning curves of the baselines and Q1 on SuperMarioBros. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
@@
Figure 11: Learning curves of the baselines and Q1 on MiniGrid-DoorKey-16×16. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
E.2 Q2
@@
Figure 12: Learning curves of the Q2 on SuperMarioBros. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
@@
Figure 13: Learning curves of the Q2 on MiniGrid-DoorKey-16×16. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
E.3 Q3
@@
Figure 14: Learning curves of the Q3 on SuperMarioBros. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
@@
Figure 15: Learning curves of the Q3 on MiniGrid-DoorKey-16×16. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
E.4 Q4
@@
Figure 16: Learning curves of the Q4 on SuperMarioBros. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
Figure 17: Learning curves of the Q4 on MiniGrid-DoorKey-16×16. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
E.5 Q5
Figure 18: Learning curves of the Q5 on SuperMarioBros. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
Figure 19: Learning curves of the Q5 on MiniGrid-DoorKey-16×16. The solid line and shaded regions represent the mean and standard deviation computed with 10 random seeds, respectively.
E.6 Q6
Figure 20: Learning curves of Q6 on Procgen-1MazeHard. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 21: Learning curves of Q6 on Procgen-AllMazeHard. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
E.7 Q7
Figure 22: Learning curves of Q7 (global+episodic exploration) on SuperMarioBros-1-1-v3 and SuperMarioBrosRandomStages-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 23: Learning curves of Q7 (global+global exploration) on SuperMarioBros-1-1-v3 and SuperMarioBrosRandomStages-v3. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
E.8 Additional Experiments for MiniGrid
Figure 24: Learning curves of four selected intrinsic rewards on eight extremely hard tasks. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 25: Learning curves on MiniGrid-MultiRoom-N2-S4-v0. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 26: Learning curves on MiniGrid-MultiRoom-N4-S5-v0. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 27: Learning curves on MiniGrid-KeyCorridorS3R3-v0. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 28: Learning curves on MiniGrid-KeyCorridorS5R3-v0. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Figure 29: Learning curves on MiniGrid-KeyCorridorS6R3-v0. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
Appendix F On-Policy RL Algorithms and Discrete Control Tasks
@@
In this section, we demonstrate the combination of RLeXplore and on-policy RL algorithms and their effectiveness on discrete control tasks. Specifically, we couple the PPO algorithm and intrinsic rewards and evaluate their performance on Montezuma Revenge, a hard exploration task from the ALE benchmark (Bellemare et al., 2013). We use the PPO implementation of CleanRL (Huang et al., 2022b) to show the adaptability of RLeXplore. Table 18 illustrates the training hyperparameters used for the experiments.
Table 18: Training hyperparameters for Montezuma Revenge.
@@
Appendix G Off-Policy RL Algorithms and Continuous Control Tasks
To showcase the generality of RLeXplore, we run additional experiments in settings different from the ones in the main paper. Concretely, we couple intrinsic rewards with soft actor-critic (SAC) (Haarnoja et al., 2018), an off-policy RL algorithm, and test their performance in Ant-UMaze, a continuous control task with sparse rewards. Table 19 illustrates the training hyperparameters used for the experiments. We show the performance of Disagreement, RND, ICM, and vanilla SAC in Figure 31. The results indicate that intrinsically-motivated agents are able to navigate the maze more efficiently, finding the goals more often than the vanilla agents that can only learn from the sparse task rewards.
We only use 3 intrinsic rewards with SAC because of the episodic nature of the other intrinsic reward methods. For example, the episodic memory in RIDE, PseudoCounts, NGU; and the episodic ellipsoid in E3B require the replay buffer to sample entire episodes instead of random rollouts. We aim to implement this logic in our RLeXplore codebase in the future.
@@ Appendix G Off-Policy RL Algorithms and Continuous Control Tasks
SAC Policy learning rate $3\cdot 10^{-4}$